OpenStudy (anonymous):

A bag contains five red marbles, seven blue marbles, and ten green marbles. one marble is drawn at random. 1. What is the probability that it will be either red or green?

OpenStudy (ddcamp):

P(a orb) = P(a) + P(b) P(Green or Red) = P(Green) + P(Red) = (10/22) + (5/22) = 15/22

OpenStudy (anonymous):

is it a permutation or combination?